I am having issues translating this word problem into an equation.
An elementary school buys tickets for a field trip to the zoo. The number of children going on the field trip is 2 times the number of adults. Tickets for children cost $4 each and tickets for adults cost $6 each. The total cost of all of the tickets is $1650, how many tickets of each type did the school purchase?
The lessons we are on has to deal with fractions but looking at this i have no idea how to make this into a fraction type equation.
The best i come up with is 2(4x)+6x = 1650. This eqaution makes x= 117.857 which i am almost poistive it is not the answer.

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

A+C=Total number of people
C=2A, so 3A is the number of people, 2/3 of whom are children.
go back to the first equation
6A+4C=1650
C=2A
6A+8A=1650
14A=1650
A=117 (6/7) which is your answer above.
Check to see
The number of children would be 235 (5/7)*$4=$942.86
Adults is 117 (6/7)*6=$707.14
Those add to $1650.
The problem did not say how many total people were there. Given the proportion of children, this is the answer.
If you try 118 adults and 236 children, it would be $1652
If you try 117 adults and 234 children, it would be $1638
There is no integer value with the C=2A constraint that works exactly for $1650.

Answer by ikleyn(52797)

.
Shortly speaking, you are right.

Your assignment problem is wrong.

Your equation is correct.

Your solution is correct, too.

They invented defective problem/data.